INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF PLATINUM FLUORIDE IN THE NEAR IN-FRARED

\underline{LEAH C. O’BRIEN}, KAITLIN WOMACK, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN, MEREDITH REDDICK, REBECCA STEIN-BERG, Department of Chemistry and Biochemistry, University of Missouri, St Louis, MO 63121-4499.

Two new bands of PtF have been recorded in the near-infrared with rotational resolution using intracavity laser absorption spectroscopy. Red-degraded bandheads are observed at 12523 and 13079 cm^{−1}, and they are identified as the (0,0) and (1,0) bands of a new electronic transition based on the observed isotopologue splitting in the bands. The results of the analysis will be presented. The gas phase PtF molecules were produced using a platinum-lined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.